The time taken to pluck the tea leaves in a certain plantation has a normal distribution with a mean of 25 hours per week and a standard deviation of 4 hours. Calculate the probability that the tea leaves can be plucked at this plantation in the following period of time:



More than 30 hours
Between 18 and 34 hours
Between 25 and 34 hours
1
Expert's answer
2019-08-28T09:24:42-0400

Let "Z=\frac{t-\mu}{\sigma}" where "t" is period of time "\mu" is mean value of t and "\sigma" is standard deviation of t.

Z has standard normal distribution

"P(a<t\leq b)=P(\frac{a-\mu}{\sigma}<\frac{t-\mu}{\sigma}\leq\frac{b-\mu}{\sigma})=\\
P(\frac{a-\mu}{\sigma}<Z\leq\frac{b-\mu}{\sigma})=P(\frac{a-25}{4}<Z\leq\frac{b-25}{4})="

"P(Z\leq\frac{b-25}{4})-P(Z\leq\frac{a-25}{4})"

1)

"P(t>30)=1-P(Z\leq\frac{30-25}{4})=\\
1-P(Z\leq\frac{5}{4})=1-0.8944=0.1056"

2)

"P(18<t<34)=\\
P(Z\leq \frac{34-25}{4}=2.25)-P(Z\leq \frac{18-25}{4}=-1.75)=\\
0.9878-0.0401=0.9477"

3)

"P(25<t<34)=\\
P(Z\leq \frac{34-25}{4}=2.25)-P(Z\leq \frac{25-25}{4}=0)=\\
0.9878-0.5=0.4878"

Answer: "P(t>30)=0.1056" , "P(18<t<34)=0.9477" , "P(25<t<34)=0.4878"